the reaction of 5 g oxygen with 5 g carbon monoxide produced 4.5 g methanol. what is the percent yield for the reaction 2H² + CO > CH³ OH
1
Expert's answer
2020-11-09T14:09:49-0500

Theoretical mass of methanol:

mt(CH3OH)=m(CO)*Mr(CH3OH)/Mr(CO)=5*32/28= 5.7 g;


Yield= ma(CH3OH)/mt(CH3OH) * 100%=4.5/5.7 *100% =78.95%

Ansewer:78.95%